OpenStudy (princezz):

How do u know when to use the product or quotient property when working with logs?

OpenStudy (anonymous):

if it's multiplication it's the sum of logs if it's division, it's the difference of logs or ln(ab) = ln(a) + ln(b) ln(a/b) = ln(a) - ln(b)

OpenStudy (ipwnbunnies):

log(a) - log(b) = log(a/b) log(a) + log(b) = log(a*b) They're just properties of logarithms.

OpenStudy (princezz):

Thanks all..... the second one makes mere sense to me :)